OpenStudy (anonymous):

Help.. Solve the system with the addition method: 3(6x+y)=-2 -3(x+5)=3y

OpenStudy (anonymous):

Simplify the left side. Change the second equation. Add it to the first.

mathslover (mathslover):

3(6x+y)=-2--------------------------1) -3(x+5)=3y--------------------------2) 1) solving 1st equation for y \[18x+3y=-2\] \[3y=-2-18x\] \[y=(-2-18x)/3\] 2) Now put the value of y in the 2nd equation on the place of y \[x+5=-y\] \[x+5=(18x+2)/3\] \[3x+15=18x+2\] \[15x=13\] \[x=13/15\] \[x+5=-y\] \[13/15+5=-y\] \[88/15=y\] \[-88/15=y\]

OpenStudy (anonymous):

18x + 3y = -2 -3x -15 = 3y Or Second Equation can be written as : -3x -3y = 15 Or 3x + 3y = -15 Form 1st equation : 3y = -2 - 18x Put this in Second Equation : 3x - 2 - 18x = -15 So, -15x = -13 So, x = 13/15 and y = -88/15...

OpenStudy (anonymous):

Thank you :) how do i do it with letters instead of numbers /:

OpenStudy (anonymous):

-a/4-b/4=1 -a/3+b/6=1